The weight of an object follows this equation:
w= Cr^-2
where c is the constant, and r is the distance that the object is from the center of the earth.
Solve the equation w= Cr^-2 for r, or in other words rearrange the equation for r.
Thank you

Answer by nyc_function(2741)

First write r^(-2) as 1/r^2
W = C/r^2
Multiply both sides by r^2.
(W)(r^2) = C
Divide both sides by W.
r^2 = C/W
Take the square root of both sides.
r = sqrt{C/W}